OpenStudy (anonymous):

A rhombus has sides of length 150 centimeters, and the angle at one of the vertices is 50°. Approximate the lengths of the diagonals to the nearest tenth of a centimeter.

OpenStudy (anonymous):



OpenStudy (anonymous):

In a rhombus, the diagonals bisect each other, so in my diagram, BE = ED. Also, in a rhombus, a diagonal bisects a vertex of the rhombus. So angle EAB = 25 degrees. Also, the diagonals are perpendicular to each other, so they form right angles.

OpenStudy (anonymous):

So we will find x, the length of BE using a trig equation: sin (25) = x/150 Cross-multiply, you get x = 150 (sin 25) x = 63.39274 So BE = ED = 63.39274, therefore, the entire diagonal BD = 2(63.39274) = 126.786

OpenStudy (anonymous):

So we found thelength of one entire diagonal. Now, you do the same for the other diagonal, AC.